OpenStudy (anonymous):

Can someone please double check my answer? What is the graph of the inequality?

OpenStudy (anonymous):

I think the correct answer is D

OpenStudy (theeric):

Hello! I agree! And, just so you know, this would be more appropriate in the "Algebra" section. But yeah! \(-3y\ge6x-3\\ y\le\dfrac{6x}{-3y}-\dfrac3{-3}=-2x+1\) So, \(y\le -2x+1\) You might even have done what I did: Look for shading on the bottom because of "less than" and look for a negative slope (down and tor the right) :)

OpenStudy (anonymous):

Thank you #theEric

OpenStudy (theeric):

You're welcome!